OpenStudy (anonymous):

solve the equation 6-9w=24 .........and i get w=2

OpenStudy (diyadiya):

Nope :D

OpenStudy (diyadiya):

6 - 9w =24 Subtract 6 from both sides -9w =24-6 -9w = 18 can you do now?

OpenStudy (diyadiya):

No i'm right :D

OpenStudy (anonymous):

w=0.5

OpenStudy (diyadiya):

How?

OpenStudy (anonymous):

sorry it is w=-2

OpenStudy (diyadiya):

Yes, Thats Right :D

OpenStudy (anonymous):

thank you

OpenStudy (diyadiya):

Anytime :)